# JOINT OPTIMIZATION OF PIECEWISE LINEAR ENSEMBLES

### v2

@@ 3.3 Multitask Dirty LASSO
We selected two multitask regression datasets from the literature: “NanoChem” [7] and “SARCOS” [24]. “NanoChem” is a group of 7 small molecule, nanoparticle, and protein datasets that were collected to evaluate multitask feature selection performance [7]. These datasets have 1,205 physico-chemical features, and the sample sizes range from 127 to 11,079. The response variables are small molecule boiling point (1), Henry’s constant (2), logP (3), melting point (4), nanoparticle logP (5), zeta potential (6), and protein solubility (7). Similarly, “SARCOS” is a 7-task dataset that models the dynamics of a robotic arm [24], with 27 features and 48,933 samples for each task. These datasets were also split into train, validation, and test sets using a 0.8/0.1/0.1 ratio.
Using JOPLEN DL with tree ensembles will provide incorrect sparsity estimates; input features will be used to create each tree (not penalized) and then to train each linear leaf (penalized). In this case, simply analyzing the leaf weights will underestimate the number of features used by the JOPLEN model. To avoid this issue, we create our partitions using random Voronoi ensembles.
All penalties were manually tuned so that all methods achieve a similar test loss for each task. Because not all datasets are equally challenging to predict, we manually tune the $\gamma^{t}$ parameter for JOPLEN DL and the equivalent for BoUTS. No equivalent parameter exists in the community implementation of DL [25].
Fig. 3: a) demonstrates that JOPLEN DL learns common and task-specific features from the SARCOS dataset. The $x$-axis represents cells grouped by task, and the $y$-axis indicates the associated input feature. Blue indicates negative weights, red indicates positive weights. b) shows regression performance, with equal performance on the diagonal. JOPLEN’s performance is equal to or greater than that of DL (orange) on most tasks, and similar to that of BoUTS (teal).
Fig. 4: Comparison of the number of common and task-specific features selected by DL, BoUTS, and JOPLEN DL for the a) SARCOS and b) NanoChem datasets (fewer features is better).
We find that JOPLEN DL and BoUTS select significantly sparser feature sets than DL does, with JOPLEN providing the sparsest sets. In Figure 3 a) we see that joint optimization enables penalties to be shared across ensemble terms, providing structured sparsity. Further, Figures 4 and 3 b) show that JOPLEN selects far fewer features than DL approach while achieving similar or superior performance on all but one task (small molecule logP (3)). This is likely because of the significant disparity in the number of features selected for this task (102 for DL vs. 6 for JOPLEN).
@@ 4 Conclusions
The JOPLEN framework enables many different penalties to be incorporated into nonlinear methods. For example, it enables a straightforward approach to feature selection in nonlinear single- and multitask settings.
We find that JOPLEN beats global refinement (GR) and significantly outperforms other state-of-the-art methods on a wide variety of real-world and synthetic datasets. We demonstrate that Laplacian; Frobenius, nuclear, $\ell_{p,1}$-norm; and Dirty LASSO regulation have straightforward extensions using JOPLEN and improve model performance. Empirically, JOPLEN shines in settings where response variables is nonlinear but has structured sparsity, such as multitask feature selection. Indeed, JOPLEN achieves similar or higher feature sparsity and performance than BoUTS (Figure 3 b)) while using suboptimal Voronoi partitions (Figure 1). Future work may use BoUTS as a partitioner for JOPLEN to improve performance while maintaining a high degree of sparsity. Combining JOPLEN with global pruning or developing new partitioning methods may also lead to further improvements. Overall, these results suggest that JOPLEN is a promising approach to improving performance on tabular datasets.
We anticipate that JOPLEN will provide improvements in general regression and classification performance across a wide variety of fields. Additionally, we expect that JOPLEN’s piecewise linear formulation will lead to improved interpretability through increased feature sparsity. Finally, because JOPLEN allows the straightforward transfer of linear penalties to the nonlinear setting, we anticipate that it will greatly simplify the implementation of nonlinear penalized optimization problems (e.g., subspace-aligned nonlinear regression via the nuclear norm, Figure 2).
@@ 5 Code availability
A JAX (CPU and GPU compatible) version of our code is available on PyPI using pip install joplen. The source code is available at https://gitlab.eecs.umich.edu/mattrmd-public/joplen-repositories/joplen. The CuPY (GPU-only) implementation used in this paper is available via the tag mlsp_2024. The training, evaluation, and plotting code for this paper is available at https://gitlab.eecs.umich.edu/mattrmd-public/joplen-repositories/joplen-tests.

### v3

@@ 3.3 Multitask Dirty LASSO
We selected two multitask regression datasets from the literature: NanoChem [7] and SARCOS [25]. NanoChem is a group of 7 small molecule, nanoparticle, and protein datasets for evaluating multitask feature selection performance [7]. These datasets have 1,205 features and 127 to 11,079 samples. The response variables are small molecule boiling point (1), Henry’s constant (2), logP (3) and melting point (4); nanoparticle logP (5) and zeta potential (6); and protein solubility (7). SARCOS is a 7-task dataset that models the dynamics of a robotic arm [25], with 27 features and 48,933 samples for each task. These datasets were also split into train, validation, and test sets using a 0.8/0.1/0.1 ratio.
Using JOPLEn DL with tree ensembles will provide biased sparsity estimates; features are used to create each tree (not penalized) and then to train each linear leaf (penalized). In this case, simply analyzing the leaf weights will underestimate the number of features used by the JOPLEn model. We avoid this issue by using random Voronoi ensembles.
All penalties were manually tuned so that all methods achieve a similar test loss for each task. Not all datasets are equally challenging, so we manually tune the $\gamma^{t}$ parameter for JOPLEn DL and BoUTS. No equivalent parameters exist in the community implementation of DL [26].
Fig. 3: a) demonstrates that JOPLEn DL learns common and task-specific features from the SARCOS dataset. The $x$-axis represents cells grouped by task, and the $y$-axis indicates the associated input feature. Blue indicates negative weights, red indicates positive weights, white is zeros. b) shows regression performance, with equal performance on the diagonal. JOPLEn’s performance is equal to or greater than that of DL (orange) on most tasks, and similar to that of BoUTS (teal).
We find that JOPLEn DL and BoUTS select significantly sparser feature sets than DL does, with JOPLEn providing the sparsest sets. Figure 3 a) shows that joint optimization shares penalties across ensemble terms, providing structured sparsity. Further, JOPLEn selects far fewer features than DL (Fig. 4) while achieving similar or superior performance on all but one task (small molecule logP (3), Fig. 3 b)). This is likely because of the significant disparity in the number of features selected for this task (102 for DL, vs. 6 for JOPLEn).
Fig. 4: The number of common and task-specific features selected by DL, BoUTS, and JOPLEn DL for the a) SARCOS and b) NanoChem datasets (fewer is better).
@@ 4 Conclusions
The JOPLEn framework enables many existing penalties for linear methods (e.g., single- and multitask feature selection) to be incorporated into nonlinear methods.
We find that JOPLEn beats global refinement (GR) and significantly outperforms other state-of-the-art methods on real-world and synthetic datasets. We demonstrate that Laplacian; Frobenius, nuclear, $\ell_{p,1}$-norm; and Dirty LASSO regulation have straightforward extensions using JOPLEn and improve model performance. Empirically, JOPLEn shines when the response variable is a nonlinear function of input features with structured sparsity, such as multitask feature selection. Such settings cannot be modeled using linear approaches such as LASSO, and JOPLEn achieves similar or higher feature sparsity and performance than BoUTS (Figure 3 b)) while using suboptimal partitions (Figure 1). Future work may use BoUTS as a partitioner for JOPLEn to improve performance while maintaining a high degree of sparsity. Combining JOPLEn with global pruning, developing new partitioning methods, and providing support for categorical features may also lead to further improvements. Overall, these results suggest that JOPLEn is a promising approach to improving performance on tabular datasets.
We anticipate that JOPLEn will improve regression and classification performance across many fields. Additionally, we expect that JOPLEn’s piecewise linear formulation will lead to improved interpretability through increased feature sparsity. Finally, because JOPLEn allows the straightforward transfer of linear penalties to the nonlinear setting, we anticipate that it will greatly simplify the implementation of nonlinear penalized optimization problems (e.g., subspace-aligned nonlinear regression via the nuclear norm, Figure 2).
@@ 5 Code availability
A JAX (CPU and GPU compatible) version of JOPLEn is available on PyPI as joplen with linked source code. The data, JOPLEn implementation, and evaluation/plotting code for this paper is available at https://gitlab.eecs.umich.edu/mattrmd-public/joplen-repositories/joplen-mlsp2024.